# Heisenberg scaling based on population coding

### v1

@@ Abstract
I Introduction
Quantum metrology is a key technology for estimating physical parameters in quantum systems with high resolution. It is known that Fisher information gives the precision limit of estimation under the $n$-copies of the unknown state is available [1, 2, 3, 4, 5, 6]. Due to the above successful role of Fisher information, many people believed that Fisher information gives the precision limit of estimation even in the quantum metrology including the unitary estimation. Therefore, many people consider that the noon state realizes Heisenberg scaling in the quantum metrology because it achieves the maximum Fisher information [7, 8, 9, 10, 11, 12, 13, 14] in the the unitary estimation. However, this maximum cannot be achieved in the sense of global estimation [15, 16, 17, 18, 19]. This fact shows that Fisher information does not reflect the global information structure. Although the references [15, 16, 17, 18, 19] pointed out this problem for Fisher information, still several papers [20, 21, 22, 23, 24, 25, 26] use Fisher information as a figure of merit for quantum metrology. To resolve this problem, it is necessary to propose an alternative figure of merit to reflect global information structure. In fact, finding such an alternative figure of merit has remained as an open problem in [17]. As an alternative figure of merit, this paper focuses on the mutual information between the parameter and the $n$-fold tensor product quantum system of our interest.
@@ I Introduction
In this paper, we study the unitary estimation with $n$-parallel application of the same unitary under several unitary models $\{U_{\theta}\}_{\theta\in\Theta}$. As the first topic, when $d$ is the number of parameters, we show that the mutual information behaves as $d\log n$, which is twice as the standard scaling $\frac{d}{2}\log n$. As the second topic, to follow the concept of population coding, we discuss how many elements in the parameter space $\Theta$ can be distinguished under $n$-parallel application. We show that the logarithm of the number of distinguishable elements is also $d\log n$ with the best initial state preparation. Another problem is the necessity of use of the ancilla or the reference system for the input state. This problem has been studied in [35, 36]. We consider this problem based under our problem setting.
Recently, the preceding studies discussed the first topic for the phase estimation [37, 38] and multi-phase estimation [39]. Also, the reference [40] studied the same topic for a generic one-parameter model. However, they did not discuss the second topic nor other models. In this paper, we study both topics under the multi-phase estimation model and the full unitary model $\mathop{\rm SU}\,(t)$.
To see the standard scaling for the mutual information, section II reviews existing result for state estimation. Section III presents our result when a general unitary representation is given. Section IV applies the above result to the case with multi-phase estimation. Section V applies the above result to the case with unitary group $\mathop{\rm SU}\,(t)$. Section VI gives the proof for a key statement in section V. Section VII makes conclusion.
II State estimation
@@
$I(A;B)=\int_{\Theta}D(\rho_{\theta}^{\otimes n}\|\rho_{A})\mu(d\theta),$ (2)
where $D(\rho\|\sigma):=\mathrm{Tr}\rho(\log\rho-\log\sigma)$.
@@
Since the number of parameters of the full model is $t^{2}-1$, (4) can be considered as a generalization of (3). When a model certain condition, using the result by [34], we can show
@@
We back to the spirit of population coding, and focus on the number $N(\{\rho^{\otimes n}_{\theta}\}_{\theta\in\Theta},\epsilon)$ of distinguishable states among $\{\rho^{\otimes n}_{\theta}\}_{\theta\in\Theta}$ with the average decoding error probability $\epsilon>0$. As shown in [41, (4.32)], using Fano inequality, we can evaluate this number as
@@
III General problem formulation
III.1 General description
@@
III.2 Mutual information
We choose an input state $\rho$ and a distribution $P$ on the system ${\cal B}=G$, where the random choice of ${\cal B}$ is denoted by the random variable $B$. We focus on the mutual information $I(Al;B)$ or $(AR;B)$, which depends on $\rho$ and $P$. To maximize the mutual information, we can restrict the input state as an input pure state $|\psi\rangle=\oplus_{\lambda\in\hat{G}_{f}}\sqrt{p_{\lambda}}|\psi_{\lambda}\rangle$ because a mixed input state can be simulated as a randomized choice of input pure states and the mutual information is convex for the randomized choice of input pure states. In this case, the mutual information is given as the von Neumann entropy of the mixture of the final states.
$I(Al;B)[P,|\psi\rangle]:=S\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)},$ (12)
where $S(\rho)$ expresses the von Neumann entropy $-\mathrm{Tr}\rho\log\rho$ of $\rho$. We consider the following optimization problem.
@@ III.2 Mutual information
$\max_{P}I(Al;B)[P,|\psi\rangle]=I(Al;B)[\mu,|\psi\rangle]=S({\cal T}_{G}(|\psi\rangle\langle\psi|)),$ (14)
@@
In particular, when $|\psi\rangle$ is the state $|\phi(p)\rangle:=\sum_{\lambda\in\hat{G}_{f}}\sqrt{p_{\lambda}}|\phi_{\lambda}\rangle$, where $p$ is a distribution on $\hat{G}_{f}$ and $|\phi_{\lambda}\rangle$ a pure state ${\cal U}_{\lambda}\otimes\mathbb{C}^{\min(d_{\lambda},ln_{\lambda})}$ such that $\mathrm{Tr}_{{\cal U}_{\lambda}}|\phi_{\lambda}\rangle\langle\phi_{\lambda}|$ is the completely mixed state on $\mathbb{C}^{\min(d_{\lambda},ln_{\lambda})}$, we have
@@
$=$ $S(\sum_{\lambda\in\hat{G}_{f}}p_{\lambda}\frac{P_{\lambda}}{d_{\lambda}\min(n_{\lambda},d_{\lambda})}$
@@
The equality holds when the input state is $|\phi(p)\rangle$ and and $p_{\lambda}=d_{\lambda}\min(ln_{\lambda},d_{\lambda})\Big{(}\sum_{\lambda^{\prime}\in\hat{G}_{f}}d_{\lambda^{\prime}}\min(ln_{\lambda^{\prime}},d_{\lambda^{\prime}})\Big{)}^{-1}$. Therefore, we have
@@
III.3 Number of distinguishable elements
Next, we consider how many elements in $G$ can be distinguished via the above process. When the input state is $\rho$, we denote the number of distinguishable states among $\{f(g)\rho f(g)^{\dagger}\}_{g\in G}$ with the average decoding error $\epsilon>0$ by $M_{Al}(\rho,\epsilon)$. We consider the following optimization problems.
@@ III.3 Number of distinguishable elements
To maximize the mutual information, we can restrict the input state as an input pure state $|\psi\rangle$ because a mixed input state can be simulated as a randomized choice of input pure states. In this case, $M_{Al}(|\psi\rangle,\epsilon)$ can be evaluated by using the Rényi entropy $S_{\alpha}\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}$ of the mixture of the final states, where $S_{\alpha}(\rho):=\frac{1}{\alpha-1}\log\mathrm{Tr}\rho^{\alpha}$ with $\alpha\neq 1$. Using the results [43, (6)] and [41, (4.67)], we have
$\log M(|\psi\rangle,\epsilon)\geq$ $\max_{P}S_{\alpha}\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}$
@@
$\log M(|\psi\rangle,\epsilon)\leq$ $\max_{P}S_{\beta}\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}$
@@
IV Multiple phase estimation
@@
where $\mathbb{N}$ is the set of non-negative integers. Since all irreducible representations are one-dimensional, $d_{\vec{n}}=1$ for $\vec{n}\in{\cal C}^{n}_{t}$ so that we do not need to attach any reference system, i.e., $R(G)_{A1}=R(G)_{AR}$. Since $|{\cal C}^{n}_{t}|=\log{n+t-1\choose t-1}$, we have
@@
The asymptotic estimation error with the above input state has been studied in [44, Sections 3 and 4]. As explained in [18, Section 4], this method is essentially the same as Kitaev’s method [45, 46], and has been implemented by [47].
@@
V $n$-tensor product representation of $\mathop{\rm SU}\,(t)$

### v2

@@ Abstract
[…]
Heisenberg scaling; quantum metrology; group symmetry; mutual information; Schur duality
1 Introduction
Quantum metrology is a key technology for estimating physical parameters in quantum systems with high resolution. It is known that Fisher information gives the precision limit of estimation under the $n$-copies of the unknown state is available [1, 2, 3, 4, 5, 6]. Due to the above successful role of Fisher information, many people believed that Fisher information gives the precision limit of estimation even in the quantum metrology including the unitary estimation. Therefore, many people consider that the noon state realizes Heisenberg scaling in the quantum metrology because it achieves the maximum Fisher information [7, 8, 9, 10, 11, 12, 13, 14] in the the unitary estimation. Although the variance-type of error is lower bounded by this maximum Fisher information, this lower bound cannot be achieved in the sense of global estimation because this lower bound is strictly larger than the minimum error under the Bayesian criterion or the minimax criterion [15, 16, 17, 18, 19, 20, 21]. This fact shows that Fisher information does not reflect the global information structure. Although the references [16, 17, 18, 19, 20] pointed out this problem for Fisher information, still several papers [22, 23, 24, 25, 26, 27, 28] use Fisher information as a figure of merit for quantum metrology. To resolve this problem, it is necessary to propose an alternative figure of merit to reflect global information structure. In fact, finding such an alternative figure of merit has remained as an open problem in [18]. As an alternative figure of merit, this paper focuses on the mutual information between the parameter and the $n$-fold tensor product quantum system of our interest.
@@ 1 Introduction
In this paper, we study the unitary estimation with $n$-parallel application of the same unitary under several unitary models $\{U_{\theta}\}_{\theta\in\Theta}$. As the first topic, when $d$ is the number of parameters, we show that the mutual information behaves as $d\log n$, which is twice as the standard scaling $\frac{d}{2}\log n$. As the second topic, to follow the concept of population coding, we discuss how many elements in the parameter space $\Theta$ can be distinguished under $n$-parallel application. We show that the logarithm of the number of distinguishable elements is also $d\log n$ with the best initial state preparation. As discussed in Sections 2 and 3.3, the number of distinguishable elements has useful meaning in estimation. Another problem is the necessity of use of the ancilla or the reference system for the input state. This problem has been studied in [37, 38]. We consider this problem based under our problem setting.
The preceding studies discussed the first topic for the phase estimation [39, 40] and multi-phase estimation [41]. Also, the reference [42] studied the same topic for a generic one-parameter model. That is, the references [39, 40, 41, 42] did not studied the full unitary model $\mathop{\rm SU}\,(t)$. The reference [43] considered the first topic for the full unitary model $\mathop{\rm SU}\,(t)$ under a different context, i.e., the contest of gate programming. However, these studies did not discuss the relation with the general framework of the representation theory because their analyses rely on their specific models. Moreover, they did not discuss the second topic. In this paper, we study both topics under the multi-phase estimation model and the full unitary model $\mathop{\rm SU}\,(t)$.
To see the standard scaling for the mutual information, section 2 reviews existing result for state estimation. Section 3 presents our result when a general unitary representation is given. Section 4 applies the above result to the case with multi-phase estimation. Section 5 applies the above result to the case with unitary group $\mathop{\rm SU}\,(t)$. Section 6 gives the proof for a key statement in section 5. Section 7 makes conclusion.
2 State estimation
@@
$I(A;B)$ $=S(\rho_{A})-\int_{\Theta}S(\rho_{\theta}^{\otimes n})\mu(d\theta),$
$=\int_{\Theta}D(\rho_{\theta}^{\otimes n}\|\rho_{A})\mu(d\theta),$ (2)
where $S(\rho)$ expresses the von Neumann entropy $-\mathrm{Tr}\rho\log\rho$ of $\rho$, and $D(\rho\|\sigma)$ expresses the quantum relative entropy $\mathrm{Tr}\rho(\log\rho-\log\sigma)$.
@@
Since the number of parameters of the full model is $t^{2}-1$, (4) can be considered as a generalization of (3). When a model satisfies a certain condition, using the result by [36], we can show
@@
We back to the spirit of population coding, and focus on the number $N(\{\rho^{\otimes n}_{\theta}\}_{\theta\in\Theta},\epsilon)$ of distinguishable states among $\{\rho^{\otimes n}_{\theta}\}_{\theta\in\Theta}$ with the average decoding error probability $\epsilon>0$. As shown in [44, (4.32)], using Fano inequality, we can evaluate this number as
@@
In estimation theory, it is also important to clarify the distance between the estimate $\theta_{guess}$ and the true value $\theta_{true}$ when we exclude a certain error probability $\epsilon>0$;
$R_{\epsilon}:=\max_{R}\{R|{\rm Pr}(\|\theta_{guess}-\theta_{true}\|>R)\leq\epsilon\},$ (7)
where $\|x\|:=\max_{j}|x_{j}|$. In the state estimation, the normalized difference $\sqrt{n}(\theta_{guess}-\theta_{true})$. asymptotically obeys a Gaussian distribution, which is called local asymptotic normality [36]. This means that the above value $R_{\epsilon}$ scales $O(n^{-1/2})$ and its coefficient is asymptotically determined by the covariance of $\sqrt{n}(\theta_{guess}-\theta_{true})$. This relation is the reason why we consider the minimum covariance in the parameter estimation in classical statistics. Since this value scales as the square of variance, we can consider that the value $R_{\epsilon}$ scales $O(n^{-1})$ in the Heisenberg scaling. That is, when we focus on $R_{\epsilon}$, the Heisenberg scaling can be considered as the case when the value $R_{\epsilon}$ scales $O(n^{-1})$. However, under the Heisenberg scaling, the variance $n(\theta_{guess}-\theta_{true})$ does not asymptotically obey a Gaussian distribution. That is, its asymptotic distribution is not fixed. Therefore, the coefficient of $R_{\epsilon}$ in the scale $O(n^{-1})$ cannot be asymptotically determined by the covariance of ${n}(\theta_{guess}-\theta_{true})$. The minimization of the variance does not have the same important role under the Heisenberg scaling as under the usual scaling.
In the following, for simplicity, we consider the one-parameter case, and assume that the parameter set is given as an interval $[a,b]$. In this case, $\lfloor(b-a)/(2R_{\epsilon})\rfloor$ elements of $[a,b]$ can be distinguished within the error probability $\epsilon$. This relation between $R_{\epsilon}$ and the number of distinguished elements can be extended to the higher dimensional case. That is, the number of distinguished elements is meaningful in the parameter estimation even under the Heisenberg scaling.
3 General problem formulation
3.1 General description
@@
3.2 Mutual information
We choose an input state $\rho$ and a distribution $P$ on the system ${\cal B}=G$, where the random choice of ${\cal B}$ is denoted by the random variable $B$. So, we have a classical-quantum state $\sum_{g\in{\cal X}}P(g)f(g)\rho f(g)^{\dagger}\otimes|g\rangle\langle g|$. We focus on the mutual information $I(Al;B)$ or $(AR;B)$, which depends on $\rho$ and $P$. When $\rho$ is decomposed as $\sum_{j}q_{j}\rho_{j}$, the mutual information is evaluated as
$I(Al;B)[P,\rho]$
::=:= $S\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})\rho f(g^{\prime})^{\dagger}\Big{)}$
$-\sum_{g\in{\cal X}}P(g)S\Big{(}f(g)\rho f(g)^{\dagger}\Big{)}$
$=$ $\sum_{g\in{\cal X}}P(g)$
$\cdot D\Big{(}f(g)\rho f(g)^{\dagger}\Big{\|}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})\rho f(g^{\prime})^{\dagger}\Big{)}$
$\leq$ $\sum_{j}q_{j}\sum_{g\in{\cal X}}P(g)$
$\cdot D\Big{(}f(g)\rho_{j}f(g)^{\dagger}\Big{\|}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})\rho_{j}f(g^{\prime})^{\dagger}\Big{)}.$ (13)
The final inequality follows from the joint convexity of relative entropy. Since any mixed state can be written as a mixture of pure states, to maximize the mutual information, we can restrict the input state as an input pure state $|\psi\rangle=\oplus_{\lambda\in\hat{G}_{f}}\sqrt{p_{\lambda}}|\psi_{\lambda}\rangle$. In this case, since the von Neumann entropy of a pure state is zero, the mutual information is given as the von Neumann entropy of the mixture of the final states.
$I(Al;B)[P,|\psi\rangle]$
$=$ $S\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}.$ (14)
We consider the following optimization problem.
@@ 3.2 Mutual information
$\max_{P}I(Al;B)[P,|\psi\rangle]=$ $I(Al;B)[\mu,|\psi\rangle]$
$=$ $S({\cal T}_{G}(|\psi\rangle\langle\psi|)),$ (16)
@@
Here, we set $|\phi_{\lambda}\rangle$ to be a pure state ${\cal U}_{\lambda}\otimes\mathbb{C}^{\min(d_{\lambda},ln_{\lambda})}$ such that $\mathrm{Tr}_{{\cal U}_{\lambda}}|\phi_{\lambda}\rangle\langle\phi_{\lambda}|$ is the completely mixed state on $\mathbb{C}^{\min(d_{\lambda},ln_{\lambda})}$. Such a state $|\phi_{\lambda}\rangle$ is called the maximally entangled state on ${\cal U}_{\lambda}\otimes\mathbb{C}^{\min(d_{\lambda},ln_{\lambda})}$.
When $|\psi\rangle$ is the state $|\phi(p)\rangle:=\sum_{\lambda\in\hat{G}_{f}}\sqrt{p_{\lambda}}|\phi_{\lambda}\rangle$, where $p$ is a distribution on $\hat{G}_{f}$, we have
@@
$=$ $S\Big{(}\sum_{\lambda\in\hat{G}_{f}}p_{\lambda}\frac{P_{\lambda}}{d_{\lambda}\min(n_{\lambda},d_{\lambda})}\Big{)}$
@@
The equality holds when the input state is $|\phi(p)\rangle$ and $p_{\lambda}=\frac{d_{\lambda}\min(ln_{\lambda},d_{\lambda})}{\sum_{\lambda^{\prime}\in\hat{G}_{f}}d_{\lambda^{\prime}}\min(ln_{\lambda^{\prime}},d_{\lambda^{\prime}})}$. Therefore, we have
@@
Overall, the optimal input state is
$\sum_{\lambda\in\hat{G}_{f}}\sqrt{\frac{d_{\lambda}\min(ln_{\lambda},d_{\lambda})}{\sum_{\lambda^{\prime}\in\hat{G}_{f}}d_{\lambda^{\prime}}\min(ln_{\lambda^{\prime}},d_{\lambda^{\prime}})}}|\phi_{\lambda}\rangle.$ (21)
@@
3.3 Number of distinguishable elements
Next, we consider how many elements in $G$ can be distinguished via the above process. When the input state is $\rho$, we denote the number of distinguishable states among $\{f(g)\rho f(g)^{\dagger}\}_{g\in G}$ with the average decoding error $\epsilon>0$ by $M_{Al}(\rho,\epsilon)$. That is,
$M_{Al}(\rho,\epsilon)$
::=:= $\max\left\{M\middle|\begin{array}[]{l}\frac{1}{M}\sum_{j=1}^{M}\mathrm{Tr}f(g_{j})\rho f(g_{j})^{\dagger}\Pi_{j}\\ \geq 1-\epsilon\end{array}\right\},$ (25)
where the maximum in (25) is taken with respect to ${M,g_{1},\ldots,g_{M}\in G,\{\Pi_{j}\}_{j=1}^{M}}$. We consider the following optimization problems.
@@ 3.3 Number of distinguishable elements
The number of distinguishable elements is related to the error radius of the estimation. We consider a distance $D$ distance on $G$ satisfying $D(g_{1},g_{2})=D(gg_{1},gg_{2})=D(g_{1}g,g_{2}g)$ for $g,g_{1},g_{2}\in G$. As the unitary version of (7), we focus on the following value.
$R_{\epsilon}:=\max_{R}\{R|{\rm Pr}(D(g_{guess},g_{true})>R)\leq\epsilon\}.$ (27)
To evaluate the above value, we introduce
$B(R):=\mu_{H}(\{g|D(g,g_{0})\leq R\}).$ (28)
Then, we have
$M_{\epsilon}(G)_{Al}B(R_{\epsilon})\leq 1$ (29)
That is, using $M_{\epsilon}(G)_{Al}$, we have the following upper bound for $R_{\epsilon}$;
$R_{\epsilon}\leq B^{-1}(M_{\epsilon}(G)_{Al}^{-1}).$ (30)
In addition, when $\{g_{j}\}$ are equally spaced, i.e., the subset $\{g_{j}\}$ forms a discrete subgroup of $G$, the maximum value (25) is achieved with a certain $\epsilon>0$. Given a discrete subgroup $\Omega:=\{g_{j}\}\subset G$, we define $B(\Omega):=\mu_{H}(\{g\in G|\min_{g}D(g^{\prime},g)=D(e,g)\})$. When $|\Omega|=M_{\epsilon}(G)_{Al}$, we have an upper bound of $R_{\epsilon}$
$R_{\epsilon}\leq R(\Omega)$
::=:= $\min_{g\in G}\{D(e,g)|\exists g^{\prime}\in\Omega,D(e,g)=D(g^{\prime},g)\}.$ (31)
However, it is not so easy to find a relation between a variance-type of error and $R_{\epsilon}$. Indeed, in the case of state estimation, the limiting distribution of the variable $\sqrt{n}(\theta_{guess}-\theta_{true})$ asymptotically obeys the Gaussian distribution, the asymptotic variance derives the asymptotic behavior of $R_{\epsilon}$. However, in the unitary estimation, the asymptotic normality does not hold [46]. Hence, we have no useful relation between a variance-type of error and $R_{\epsilon}$.
To maximize the mutual information, we can restrict the input state as an input pure state $|\psi\rangle$ because a mixed input state can be simulated as a randomized choice of input pure states. In this case, $M_{Al}(|\psi\rangle,\epsilon)$ can be evaluated by using the Rényi entropy $S_{\alpha}\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}$ of the mixture of the final states, where $S_{\alpha}(\rho):=\frac{1}{\alpha-1}\log\mathrm{Tr}\rho^{\alpha}$ with $\alpha\neq 1$. Using the results [47, (6)] and [44, (4.67)], we have
$\log M(|\psi\rangle,\epsilon)$
$\geq$ $\max_{P}S_{\alpha}\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}$
@@
$\log M(|\psi\rangle,\epsilon)$
$\leq$ $\max_{P}S_{\beta}\Big{(}\sum_{g^{\prime}\in{\cal X}}P(g^{\prime})f(g^{\prime})|\psi\rangle\langle\psi|f(g^{\prime})^{\dagger}\Big{)}$
@@
4 Multiple phase estimation
@@
where $\mathbb{N}$ is the set of non-negative integers. Since all irreducible representations are one-dimensional, $d_{\vec{n}}=1$ for $\vec{n}\in{\cal C}^{n}_{t}$ so that we do not need to attach any reference system, i.e., $R(G)_{A1}=R(G)_{AR}$. Since $|{\cal C}^{n}_{t}|={n+t-1\choose t-1}$, we have
@@
The asymptotic estimation error with the above input state has been studied in [46, Sections 3 and 4]. As explained in [19, Section 4], this method is essentially the same as Kitaev’s method [48, 49], and has been implemented by [50].
@@
However, when we focus on the error $\sin^{2}\frac{\theta_{guess}-\theta_{true}}{2}$, the input state (39) does not coincide with the optimal input [46, Sections 3 and 4], [51, Section 10.2]. Further, the input state (39) does not achieve Heisenberg limit in this sense [46, Sections 3 and 4], [52]. However, the references [46, Sections 3 and 4] [19, Section 4] show that the input state (39) achieves the Heisenberg scaling in the sense limiting distribution.
5 $n$-tensor product representation of $\mathop{\rm SU}\,(t)$